OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

When a number is a solution to an equation, it means if you plug the number in to that equation the equation makes sense. Equations usually have an equal sign, so just complete both sides of the equation and see if they equal each other. An example: 5+x = 7 so when x is 2, 5+x = 5+2 = 7 therefore 2 is the solution to the equation.

OpenStudy (anonymous):

thank you